If a chord of the parabola y^2 = 4ax is a normal at one of its ends, show that its mid-point lies on the curve 2(x−2a) = (y^2 /a) +( 8a^3/ y^2) . Prove that the shortest length of such a chord is 6a√3
1
Expert's answer
2020-06-30T18:25:11-0400

If the normal at the point "P(at_1^2, 2at_1)" meets the parabola "y^2=4ax" again at "Q(at_2^2, 2at_2)" then


"t_2=-t_1-{2\over t_1}"

"Q=(a(t_1+{2\over t_1})^2, -2a(t_1+{2\over t_1}))"

Let "M(x,y)" be the midpoint of "PQ"


"(x, y)=\big(\dfrac{at_1^2+a(t_1+\dfrac{2}{t_1})^2}{2}, \dfrac{2at_1-2a(t_1+\dfrac{2}{t_1})}{2}\big)"

"t_1=-\dfrac{2a}{y}"

"x=at_1^2+2a+\dfrac{2a}{t_1^2}"

"x-2a=\dfrac{4a^3}{y^2}+\dfrac{y^2}{2a}"


"2(x-2a)=\dfrac{y^2}{a}+\dfrac{8a^3}{y^2}"

"PQ^2=(-at_1^2+a(t_1+\dfrac{2}{t_1})^2)^2+(-2a(t_1+{2\over t_1})-2at_1)^2"

"PQ^2=(4a+\dfrac{4a}{t_1^2})^2+(4at_1+\dfrac{4a}{t_1})^2"

"PQ^2=16a^2(1+\dfrac{2}{t_1^2}+\dfrac{1}{t_1^2}+t_1^2+2+\dfrac{1}{t_1^4})"

"PQ^2=16a^2(t_1^2+\dfrac{3}{t_1^2}+\dfrac{1}{t_1^4}+3)"

Consider the function

"f(u)=u+\dfrac{3}{u}+\dfrac{1}{u^2}+3, u>0"

"f'(u)=1-\dfrac{3}{u^2}-\dfrac{2}{u^3}"

"f'(u)=0=>\dfrac{u^3-3u-2}{u^3}=0"

"u^2(u-2)+2u(u-2)+(u-2)=0"

"(u-2)(u^2+2u+1)=0"

"(u-2)(u+1)^2=0"

Since "u>0," we take "u=2"

"0<u<2, f'(u)<0, f(u) \ decreases"

"u>2, f'(u)>0, f(u) \ increases"

"f(2)=2+\dfrac{3}{2}+\dfrac{1}{4}+3=\dfrac{27}{4}"

"PQ^2\geq16a^2(\dfrac{27}{4})"

"PQ^2\geq108a^2"


"PQ\geq6\sqrt{3}a, a>0"